OpenStudy (anonymous):

What is 6 2/3 divided by 3 3/4?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

I have to show workkkkk

OpenStudy (anonymous):

I got 18 8/9

OpenStudy (anonymous):

thanks for volunteering my services without even askin @e.cociuba!

OpenStudy (anonymous):

but i THINK i can figure it out..........................

OpenStudy (anonymous):

yes. That's what I did

OpenStudy (anonymous):

i was half way thru my drawing when u finished... haha but now we are just littering this poor gals post up with comments... anyhew gotta run, love ta hang... you know... any brb!

OpenStudy (anonymous):

Sort of but how did you get 15/1? Isn't it 15/4?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

omg lol

OpenStudy (anonymous):

\[ \frac{ 6 \frac{ 2 }{ 3 } }{ 3\frac{ 3 }{ 4 } }\]

OpenStudy (anonymous):

so thats ur problem/question?

OpenStudy (anonymous):

Yeah, I got 5 1/3

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

? How?

OpenStudy (e.cociuba):

Ya I have no clue where you keep coming up w/ these answers. @some_someone Don't u use the stay, change, flip method?

OpenStudy (anonymous):

I did use the flip method.

OpenStudy (anonymous):

It's 1 7/9

OpenStudy (anonymous):

@elenabrooke it would be "easier" to compute and work it out if you made the fractions to improper fractions. \[6\frac{ 2 }{ 3 } = \frac{ 20 }{ 3 }\] \[3\frac{ 3 }{ 4 } = \frac{ 15 }{ 4 }\]

OpenStudy (e.cociuba):